OpenStudy (anonymous):

Which equation is a quadratic equation? 5x − 2 = y y + 15x2 = (3x − 2)(5x + 7) y = 3x(5 − x) + 7 y + 2 = (3x2 + 10)(x − 1)

OpenStudy (anonymous):

quadratic equation is any equation having x^2 and is equal to zero

OpenStudy (anonymous):

HINT: Quadratic equations have X^2 in it

OpenStudy (anonymous):

So is it y + 15x2 = (3x − 2)(5x + 7)

OpenStudy (anonymous):

there are more terms with the highest power of 2 to x (x^2) please multiply all out and see which one equals zero

OpenStudy (anonymous):

I know there is more then one.

OpenStudy (anonymous):

I am asking you if y + 15x2 = (3x − 2)(5x + 7) is the right one.

OpenStudy (anonymous):

How did u get to this answer

OpenStudy (anonymous):

So its not the answer?

OpenStudy (amistre64):

it looks right, but it has a flaw in it .....

OpenStudy (amistre64):

when you multiply that out on the right side; you get a 3x*5x = 15x^2 right?

OpenStudy (amistre64):

y + b = b + c ; when we simplify - b -b ------------- y = c ; the term that is alike on each side goes away

OpenStudy (anonymous):

okay thanks

OpenStudy (amistre64):

yw, so what "looks" like it could be a quadratic in that case is just a disguised linear :/ sneaky wasnt they :)